The manufacturer sells an article in lots of hundred and agree to pay the purchaser a penalty of no if five or more articles in a lot are defective. He knows from past experience that, on average, 2% of the articles are defective. It costs €2 to make each article, what should the selling price be if the manufacturers over all profit is to be 25%?


1
Expert's answer
2021-05-20T17:18:41-0400

"\hat{p}=\dfrac{5}{100}=0.05"


On an average 2% of the article is defective-

"p=0.02"

Number of defective items "= 0.02\times 100=2"


Cost of the manufacturing of items "= 100\times 2=200"



Profit "= \dfrac{Sp-Cp}{Cp}\times 100"


"\Rightarrow25=\dfrac{Sp-200}{200}\times 100\\[9pt]\Rightarrow Sp-200=50\Rightarrow Sp=250"


Hence The selling price will be 250 .